OpenStudy (anonymous):

Admission to the museum is $12 for children and $15 for adults. On a certain day, 2450 people go to the museum and $33,456 is collected. How many children and how many adults went to the museum? children adults

OpenStudy (anonymous):

i love math :D

OpenStudy (anonymous):

c = number of children ---- a = number o adults c + a = 2450 --- a = 2450 - c 12c + 15a = 33456 12c + 15(2450 - c) = 33456 -3c = 33456 - 15(2450) = -3294 c = 1098 ----- a = 1352 << answer

OpenStudy (anonymous):

dose that help?

OpenStudy (anonymous):

im a god at math :P

OpenStudy (anonymous):

thank you so much :D

OpenStudy (anonymous):

np